OpenStudy (anonymous):

• When three factors are multiplied, how is the sign of the product determined?

OpenStudy (anonymous):

For example 57 . Look : 5+7=12___>1+2=3/3=1 Got it ?

OpenStudy (anonymous):

+ all of the numbrs have and / on 3 . Just this :)

OpenStudy (anonymous):

sorry I don't get what you just did

OpenStudy (tkhunny):

Count the negative values. Odd? Negative. Even? Positive.

OpenStudy (anonymous):

ok